News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 9 December 2005 9.00am EST

During the last week Chandra completed the observing schedule as planned.

Of note this week was the decision to increase the projected maximum temperature allowed in mission planing of the EPHIN radiation detector from 96F to 110F. The new limit was adopted to relax increasingly tight mission planning thermal constraints at normal sun attitudes. The new limit will result in an increased number of EPHIN 27 volt rail drops, however following review, this has been assessed as acceptable by the engineering team and EPHIN PI.

A Chandra press release was issued on Dec 1 describing observations of the Perseus galaxy cluster that show evidence of energetic plumes extending 300,000 light years into the massive cluster. The plumes, due to explosive venting from the vicinity of a supermassive black hole, provide new evidence of the influence a black hole over intergalactic distances. For further details see: http://chandra.harvard.edu/press/05_releases/press_120105.html

The schedule of targets for next week is shown below.

------------------------------------------
SDSSJ0812+4018          ACIS-S       Dec 11
SN 2005gj               ACIS-S
NGC 5253                ACIS-S       Dec 12
15 Mon                  ACIS-S/HETG
Radiation Belts                      Dec 13
RX J0720.4-3125         HRC-S/LETG
DLS J1359-1055          ACIS-I       Dec 14
1RXS J205334.0-063      ACIS-S
Q2130-431               ACIS-S
IRAS05436-0007          ACIS-I
NGC6872                 ACIS-I
SDSS J210757.67-06      ACIS-S       Dec 15
DG Tau                  ACIS-S
RX J0720.4-3125         HRC-S/LETG
Radiation Belts
RX J0720.4-3125         HRC-S/LETG   Dec 16
NGC6872                 ACIS-I
DG Tau                  ACIS-S       Dec 17
IRAS03359+1523          ACIS-S
RX J0720.4-3125         HRC-S/LETG
NGC 2500                ACIS-S
Q2131-437               ACIS-S
Q2122-444               ACIS-S
NGC 7090                ACIS-S       Dec 18

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports